# SLICE: SLO-Driven Scheduling for LLM Inference on Edge Computing Devices

### v1

@@ I Introduction
(a) Decode latency varing with batch size
(b) Token throughput varing with batch size
@@
In the second phase (i.e., the rate allocation phase), the decoding rate is dynamically adjusted based on the SLO requirements of each task. To minimize system overhead, this paper proposes an innovative decode-mask matrix that enables cycle-based scheduling to achieve rate control for individual tasks. Specifically, the number of columns in the matrix represents the total number of decoding operations per cycle, while each row corresponds to one scheduled task, recording its decoding schedule within the cycle. In the matrix, setting the $n$th column of a row to ”1” indicates that the corresponding task participates in the $n$th decoding operation of that cycle. For tasks with stricter TPOT SLO requirements (i.e., those requiring a higher decoding rate), the ”1”s are distributed more densely across their respective rows, thereby enabling a higher decoding rate. During the actual scheduling process, the system performs column-wise scanning starting from the first column in each cycle, triggering a decoding operation upon completing the scan of each column. For every scan, the system groups all tasks corresponding to rows marked ”1” in the current column into a single decoding batch. Through this iterative scanning and dynamic batch aggregation mechanism, the system can dynamically adjust the decoding rate of each task according to its SLO requirements.
@@
  • This paper investigates the distinctive challenges in LLM task scheduling for edge computing scenarios, where different tasks exhibit significant variations in SLOs for inference performance;
  • The paper proposes a utility-maximizing task selection mechanism that optimizes task scheduling upon arrival, achieving dual objectives of real-time task deadline assurance and maximum overall SLO attainment;
  • The paper proposes a decode-mask matrix periodic scanning mechanism that accurately controls the decoding rate of scheduled tasks even under various load conditions;
@@
The rest of this paper is organized as follows. Section II presents the research background and related works. Section III and IV provide the motivations and design of SLICE, respectively. Section VI presents the validation of SLICE, and Section VII concludes the research contributions.
II Background and Related Works
@@ IV SLICE: SLO-Driven Schduling
To address the aforementioned challenges, this chapter proposes a systematic solution. Section IV-A constructs a utility maximization model to transform the dual-objective optimization scheduling problem into a mathematical optimization problem, identifying the optimal trade-off between real-time guarantees and overall SLO attainment optimization by maximizing system-wide comprehensive utilities. Section IV-B introduces an offline SLICE scheduling strategy that systematically addresses the aforementioned optimization challenges through its two core components—task scheduling selection and task rate allocation. Section IV-C provides a task selection scheme based on a benefit-rate-priority decision framework, enabling maximizing system-wide comprehensive benefits to resolve Challenge 1. Section IV-D innovatively proposes an SLO-aware decoding mask matrix to allocate the decoding rates of scheduled tasks, tackling Challenge 2. Section IV-E further realizes the effective transformation of the offline SLICE into an online strategy, ensuring efficient and stable execution in real-time scenarios.
@@ IV-D Rate Allocation
Step 2: SLO-aware decoding (line 12-33). During decoding execution, a column-wise scanning mechanism is adopted. The system starts scanning from the first column of the matrix sequentially, and advances to the next column after each decoding step. During the scanning of any column, all tasks corresponding to the elements with a value of 1 in that column are grouped into a batch for decoding (line 16-19). When the scanning of the last column is completed, a full scheduling cycle is finished, and the system automatically resets to the first column to begin a new cycle. Through this cyclic column-scanning mechanism, the system can accurately achieve dynamic allocation of differentiated rates for each task. The algorithm will continue execution until either all tasks are completed or an interruption event occurs.
As shown in Fig. 4, during the decoding execution phase, the SLICE algorithm performs periodic scheduling through column-wise scanning: each scheduling cycle scans sequentially from the 0th column to the 6th column from left to right, after which the next cycle restarts scanning from the 0th column and repeats cyclically. During each scan, the system groups tasks corresponding to the positions with a value of 1 in the current column into a batch for decoding. For example, when scanning the 2nd column, if the 0th and 1st elements in that column are 1, the SLICE algorithm will group task0 and task1 into a batch for decoding.
@@ VI-A Experimental Settings
    Real-time tasks (20 tokens/s), such as machine control and navigation planning [10], demand strict adherence to response rates to ensure critical tasks are completed within deadlines (i.e., 1.5s).
@@
In the workload design, task arrivals are simulated using a Poisson process, with eight incrementally increasing request arrival rates (ranging from 0.1 to 7.0 tasks/sec). This approach systematically examines the system’s performance evolution from idle to overloaded states while assessing SLO compliance across different task types. Such a refined experimental setup enables a thorough analysis of the system’s performance characteristics and potential bottlenecks under varying load conditions.

### v2

@@ I Introduction
(a) Decode latency VS. batch size
(b) Token throughput VS. batch size
@@
In the second phase (i.e., the rate allocation phase), the decoding rate is dynamically adjusted based on the SLO requirements of each task. To minimize system overhead, this paper proposes an innovative decode-mask matrix that enables cycle-based scheduling to achieve rate control for individual tasks. Specifically, the number of columns in the matrix represents the total number of decoding operations per cycle, while each row corresponds to one scheduled task, recording its decoding schedule within the cycle. In the matrix, setting the $n$th column of a row to ”1” indicates that the corresponding task participates in the $n$th decoding operation of that cycle. For tasks with stricter TPOT SLO requirements (i.e., those requiring a higher decoding rate), the ”1”s are distributed more densely across their respective rows, thereby enabling a higher decoding rate. During the actual scheduling process, the system performs column-wise scanning starting from the left to right column in each cycle, triggering a decoding operation upon completing the scan of a column. For every scan, the system groups all tasks corresponding to rows marked ”1” in the current column into a single decoding batch. Through this iterative scanning and dynamic batch aggregation mechanism, the system can dynamically adjust the decoding rate of each task according to its SLO requirements.
@@
  • This paper investigates the distinctive challenges in LLM task scheduling for edge computing scenarios, where different tasks exhibit significant variations in SLOs for inference performance.
  • The paper proposes a utility-maximizing task selection mechanism that optimizes task scheduling upon arrival, achieving dual objectives of real-time task deadline assurance and maximum overall SLO attainment.
  • The paper proposes a decode-mask matrix periodic scanning mechanism that accurately controls the decoding rate of scheduled tasks even under various load conditions.
@@
The rest of this paper is organized as follows. Section II presents the research background and related work. Section III and IV provide the motivations and design of SLICE, respectively. Section VI presents the validation of SLICE, and Section VII concludes the research contributions.
II Background and Related Work
@@ IV SLICE: SLO-Driven Schduling
To address the aforementioned challenges, this chapter proposes a systematic solution. Section IV-A constructs a utility maximization model to transform the dual-objective optimization scheduling problem into a mathematical optimization problem, identifying the optimal trade-off between real-time guarantees and overall SLO attainment optimization by maximizing system-wide comprehensive utilities. Section IV-B introduces an offline SLICE scheduling strategy that systematically addresses the aforementioned optimization challenges through its two core components—task scheduling selection and task rate allocation. Section IV-C provides a task selection scheme based on a utility-rate-priority decision framework, enabling maximizing system-wide comprehensive utility to resolve Challenge 1. Section IV-D innovatively proposes an SLO-aware decoding mask matrix to allocate the decoding rates of scheduled tasks, tackling Challenge 2. Section IV-E further realizes the effective transformation of the offline SLICE into an online strategy, ensuring efficient and stable execution in real-time scenarios.
@@ IV-D Rate Allocation
Step 2: SLO-aware decoding (line 12-33). During decoding execution, a column-wise scanning mechanism is adopted. The system sequentially scans the matrix columns from left to right, moving to the next column after completing the scanning and decoding of the current column. During the scanning of a column, all tasks corresponding to the elements with a value of 1 in that column are grouped into a batch for executing a decoding step (line 16-19). When the scanning and decoding of the rightmost column is completed, a full scheduling cycle is finished, and the system automatically resets to the leftmost column to begin a new cycle. Through this cyclic column-scanning mechanism, the system can accurately achieve dynamic allocation of differentiated rates for each task. The algorithm will continue execution until either all tasks are completed or an interruption event occurs.
As shown in Fig. 4, during the decoding execution phase, the SLICE algorithm performs periodic scheduling through column-wise scanning. Each scheduling cycle scans sequentially from the 0th column to the 5th column, after which the next cycle restarts scanning from the 0th column and repeats cyclically. During each scan, the system groups tasks corresponding to the positions with a value of 1 in the current column into a batch for a decoding step. For example, when scanning the 2nd column, if the 0th and 1st elements in that column are 1, the SLICE algorithm will group task0 and task1 into a batch and perform one decoding step.
@@ VI-A Experimental Settings
    Real-time tasks, such as machine control and navigation planning [10], demand strict adherence to response rates (i.e., at least 20 tokens/s) to ensure critical tasks are completed within deadlines (i.e., 1.5s).
@@
In the workload design, task arrivals are simulated using a Poisson process, with ten incrementally increasing request arrival rates (ranging from 0.1 to 7.0 tasks/sec). This approach systematically examines the system’s performance evolution from idle to overloaded states while assessing SLO compliance across different task types. Such a refined experimental setup enables a thorough analysis of the system’s performance characteristics and potential bottlenecks under varying load conditions.